If we tripled an ohmic wire RADIUS, then its resistance..?
Answer #1

The electrical resistance of a wire is given by

R = (p * L)/A

where,
R = Resistance of wire
p = Ressistivity of wire
L = length of wire
A = area of cross section = 3.14 * (r)^2
where r = radius of wire.

If we triple radius,
Anew = 3.14 * (3*r)^2

Anew = 3.14 * r^2 * 9
Anew = 9 * A

Rnew = (p * L)/Anew
Rnew = (p * L)/(A * 9)
Rnew = R/9

So we see, Resistance of the wire will reduce by the factor of 9.